# Observation of Cosmic-Ray Anisotropy in the Southern Hemisphere with Twelve Years of Data Collected by the IceCube Neutrino Observatory

### v1

@@ 3.6 Angular Scale Decomposition
Figure 8: Relative intensity (top) and statistical significance (bottom) of the residual sky maps at low- (left) and high-energy (right) events ($E<18$ TeV and $>320$ TeV, respectively). The small-scale structure is made visible by subtracting the best-fit dipole and quadrupole components from the large-scale structure. All maps are shown in J2000 equatorial coordinates and use an angular smoothing radius of $5^{\circ}$ for the low-energy sample and $20^{\circ}$ for the high-energy one. The dashed line and triangle mark the galactic plane and center.
@@ 4.1 Interference with Earth’s Orbital Motion
As the Earth revolves around the Sun, we observe an apparent excess of cosmic rays in the direction of the Earth’s motion and a deficit in the opposite direction. Such an anisotropy, referred to as the Compton–Getting Effect (Compton & Getting, 1935; Gleeson & Axford, 1968) , explains the directional dependence caused by the movement of an observer with velocity $v$ through a flux of particles or photons. This effect can be observed in a solar time frame where the Sun remains at a fixed position and the Earth completes one revolution in 365.242 days. The resulting pattern of arrival directions is referred to as the “solar dipole” and is a well-understood phenomenon used to verify the reliability and stability of an anisotropy analysis. We studied this effect extensively and we documented it in various publications such as Abbasi et al. (2011, 2012b); Aartsen et al. (2016a).
@@
The interference between anisotropies originating in different time frames produces overlapping frequency sidebands. The sidereal day is 3.93 minutes shorter than a solar day; to assess the impact of the orbital interference on the celestial signal, we can look at the anti-sidereal time frame, where a day is 3.93 minutes longer than a solar day. The observation of a statistically significant antisidereal anisotropy can be used to either apply a correction to recover the actual anisotropy (with a method developed by Farley & Storey (1954) and used by Guillian et al. (2007)) or estimate a systematic error, as done in Aartsen et al. (2016a).
In this work, we produced cosmic-ray relative intensity sky maps in the anti-sidereal time frame, calculated the one-dimensional projection of relative intensity in the RA coordinates (see Sec. 3.4), and used it to determine the RMS spread of the relative intensity at each RA bin for each year of our data sample. Figure 9 shows that the systematic errors estimated with this method are of the same order as the statistical errors, a major improvement over the previous results published in Aartsen et al. (2016a). The dramatic reduction of the orbital interference systematic uncertainty is due to the data being split into calendar years. In past analyses, where we used data collected during IceCube’s construction, we had to rely on the experiment’s configuration years, which were anywhere from 8 to 14 months long. Another reason for the small orbital interference systematic uncertainties stems from an absence of large data gaps due to the high duty cycle of the IceCube observatory over the last twelve years (see Table 1). Histograms of time gaps between events showed no notable peaks beyond $10^{-2}$ seconds. Since the angular rotation of the Earth in $10^{-2}$ s (($4.2\times 10^{-5})^{\circ}$) is much less than the pixel size ($(0.84^{\circ})^{2}$), any peaks on this scale and smaller should have no significant effect on our directional reconstruction.
@@ 4.2 Anisotropy Stability Over Time
We investigated whether the cosmic-ray anisotropy is constant in time across the twelve years considered in the analysis. To do so, we looked at the one-dimensional projection of the relative intensity — binned in RA — for each calendar year. Figure 9 shows the residual between each year’s relative intensity from that of the entire data sample. The solid error bars at each point are statistical uncertainties. The shaded boxes around each data point represent systematic uncertainties calculated from the anti-sidereal distribution for each year (see Sec. 4.1). The effect of switching to calendar years is seen in the reduced size of the systematic uncertainties compared to previous analyses (Aartsen et al., 2016a) — they are now comparable in size to statistical uncertainties. The black band at each RA bin corresponds to the relative intensity from the entire twelve-year sample.
@@ 5 Summary
We often use the dipole component of the observed anisotropy as a proxy for cosmic-ray diffusive models in the ISM. However, the abovementioned studies highlight the need to carefully measure the cosmic-ray arrival direction distribution at small scales (i.e., with $-\ell$ multipole moments) as the angular power spectrum reflects the turbulence present in interstellar space. To determine the dipole component, we updated our measurement by fitting a two-dimensional arrival direction distribution sky map that takes into account the limited FoV of the IceCube in-ice array (see Sec. 3.4). In Fig. 4, we have presented the amplitude and phase of the dipole component and compared our results with those of other experiments. We have also presented the angular pseudo-power spectrum to show the energy dependence of the different multipole modes and highlighted the importance of full-sky cosmic-ray anisotropy observations. It is important to note that all reported observations are only sensitive to the horizontal components of the anisotropy, as the data-driven estimation of the reference maps absorbs the equatorial north-south component.
@@ 6 Outlook
Future advancements in instrumentation, such as IceCube-Gen2 and the Southern Wide-field Gamma-ray Observatory (SWGO), are expected to enhance our observations even further. IceCube-Gen2 is designed to have a broader aperture, which will enable surface-to-in-ice coincidence studies, consequently improving energy and mass sensitivity in the PeV energy range (Aartsen et al., 2021; Schroeder et al., 2023; IceCube-Gen2, 2023). In addition, the SWGO observatory in the Southern Hemisphere will provide modern instrumentation with improved energy and mass composition sensitivity and significantly contribute to cosmic-ray physics, including anisotropy measurements (Taylor, 2021).

### v2

@@ 3.6 Angular Scale Decomposition
Figure 8: Relative intensity (top) and statistical significance (bottom) of the residual sky maps at low- (left) and high-energy (right) events ($E<18$ TeV and >320>320> 320 TeV, respectively). The small-scale structure is made visible by subtracting the best-fit dipole and quadrupole components from the large-scale structure. All maps are shown in J2000 equatorial coordinates and use an angular smoothing radius of $5^{\circ}$ for the low-energy sample and $20^{\circ}$ for the high-energy one. The dashed line and triangle mark the galactic plane and center.
@@ 4.1 Interference with Earth’s Orbital Motion
As the Earth revolves around the Sun, we observe an apparent excess of cosmic rays in the direction of the Earth’s motion and a deficit in the opposite direction. Such an anisotropy, referred to as the Compton–Getting Effect (Compton & Getting, 1935; Gleeson & Axford, 1968) , explains the directional dependence caused by the movement of an observer with velocity $v$ through a flux of particles or photons. This effect can be observed in a solar time frame where the Sun remains at a fixed position and the Earth completes one revolution in 365.242 days. The resulting pattern of arrival directions is referred to as the “solar dipole” and is a well-understood phenomenon used to verify the reliability and stability of an anisotropy analysis. We studied this effect extensively and we documented it in various publications such as Abbasi et al. (2011, 2012a); Aartsen, M. G. et al. (2016a).
@@
The interference between anisotropies originating in different time frames produces overlapping frequency sidebands. The sidereal day is 3.93 minutes shorter than a solar day; to assess the impact of the orbital interference on the celestial signal, we can look at the anti-sidereal time frame, where a day is 3.93 minutes longer than a solar day. The observation of a statistically significant antisidereal anisotropy can be used to either apply a correction to recover the actual anisotropy (with a method developed by Farley & Storey (1954) and used by Guillian et al. (2007)) or estimate a systematic error, as done in Aartsen, M. G. et al. (2016a).
In this work, we produced cosmic-ray relative intensity sky maps in the anti-sidereal time frame, calculated the one-dimensional projection of relative intensity in the RA coordinates (see Sec. 3.4), and used it to determine the RMS spread of the relative intensity at each RA bin for each year of our data sample. Figure 9 shows that the systematic errors estimated with this method are of the same order as the statistical errors, a major improvement over the previous results published in Aartsen, M. G. et al. (2016a). The dramatic reduction of the orbital interference systematic uncertainty is due to the data being split into calendar years. In past analyses, where we used data collected during IceCube’s construction, we had to rely on the experiment’s configuration years, which were anywhere from 8 to 14 months long. Another reason for the small orbital interference systematic uncertainties stems from an absence of large data gaps due to the high duty cycle of the IceCube observatory over the last twelve years (see Table 1). Histograms of time gaps between events showed no notable peaks beyond $10^{-2}$ seconds. Since the angular rotation of the Earth in $10^{-2}$ s (($4.2\times 10^{-5})^{\circ}$) is much less than the pixel size ($(0.84^{\circ})^{2}$), any peaks on this scale and smaller should have no significant effect on our directional reconstruction.
@@ 4.2 Anisotropy Stability Over Time
We investigated whether the cosmic-ray anisotropy is constant in time across the twelve years considered in the analysis. To do so, we looked at the one-dimensional projection of the relative intensity — binned in RA — for each calendar year. Figure 9 shows the residual between each year’s relative intensity from that of the entire data sample. The solid error bars at each point are statistical uncertainties. The shaded boxes around each data point represent systematic uncertainties calculated from the anti-sidereal distribution for each year (see Sec. 4.1). The effect of switching to calendar years is seen in the reduced size of the systematic uncertainties compared to previous analyses (Aartsen, M. G. et al., 2016a) — they are now comparable in size to statistical uncertainties. The black band at each RA bin corresponds to the relative intensity from the entire twelve-year sample.
@@ 6 Outlook
Future advancements in instrumentation, such as IceCube-Gen2 and the Southern Wide-field Gamma-ray Observatory (SWGO), are expected to enhance our observations even further. IceCube-Gen2 is designed to have a broader aperture, which will enable surface-to-in-ice coincidence studies, consequently improving energy and mass sensitivity in the PeV energy range (Aartsen, M. G. et al., 2021; Schroeder et al., 2023; IceCube-Gen2, 2023). In addition, the SWGO observatory in the Southern Hemisphere will provide modern instrumentation with improved energy and mass composition sensitivity and significantly contribute to cosmic-ray physics, including anisotropy measurements (Taylor, 2021).